OpenStudy (anonymous):

p + 3w= −94 −2p + W= −22 (p, W) =

pooja195 (pooja195):

Im guessing we are using the addition method here since the first ones are opposites they will cancel out

pooja195 (pooja195):

\[3w+w\]

OpenStudy (sleepyjess):

p-2p is -p

pooja195 (pooja195):

huh????

OpenStudy (sleepyjess):

you said that p cancels..

OpenStudy (anonymous):

www.mathway.com

Nnesha (nnesha):

@bani2100 still need help ???

OpenStudy (anonymous):

you can use elimination Multiply the first equation with 2 2(p +3 w) =2(-94) 2p + 6w = -188 (1) Now add (1) to your second equation to eliminate p 2p + 6w = -188 -2p + w = -22 ----------------- 7w = - 210 w = -30 Substitute w to find p 2p + 6(-30) = -188 2p - 180 = -188 2p = -8 p= -4

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

Yes thank you!